Fairmount Travel Gear produces backpacks and sells them to vendors who sell them under their own label. The cost of one of its backpacks follows.

Materials $ 18.50
Labor 12.50
Variable overhead 5.50
Fixed overhead ($2,957,500 per year; 455,000 units per year) 6.50
Total $ 43.00

Riverside Discount Mart, a chain of low-price stores, has asked Fairmount to supply it with 24,000 backpacks for a special promotion Riverside is planning. Riverside has offered to pay Fairmount a unit price of $47 per pack. The regular selling price is $65. The special order would require some modification to the basic model. These modifications would add $4.50 per unit in material cost, $2.00 per unit in labor cost, and $1.00 in variable overhead cost. Although Fairmount has the capacity to produce the 24,000 units without affecting its regular production of 455,000 units, a one-time rental of special testing equipment to meet Riverside’s requirements would be needed. The equipment rental would be $87,000 and would allow Fairmount to test up to 55,000 units.

Required:

a. Prepare a schedule to show the impact of filling the Riverside order on Fairmont’s profits for the year.

b. Do you agree with the decision to accept the special order?

c. Considering only profit, determine the minimum quantity of backpacks in the special order that would make it profitable.

(All Costs in Thousands of Dollars)
Status Quo 455,000 Units Alternative 479,000 Units Difference
Sales revenue
Less variable costs:
Materials
Labor
Variable overhead
Total variable cost $0.0 $0.0
Contribution margin
Less: Fixed costs
Operating profit (loss)
Answer #1

Solution :

(1)

Status Quo 455,000 Units (In '000) Alternativ 479,000 Units (In '000) Difference (In '000)
Sales Revenue $ 29,575 $ 30,703 $ 1,128 Higher
Less : Variable Costs
Direct Material $ 8,417.5 $ 8,969.50 $ 552 Higher
Direct Labor $ 5,687.5 $ 6,035.5 $ 348 Higher
Variable OH $ 2,502.5 $ 2,658.50 $ 156 Higher
Total Variable Cost $ 16,607.5 $ 17,663.5 $ 1056 Higher
Contribution Margin $ 12,967.5 $ 13,039.5 $ 72 Higher
Less : Fixed OH $ 2,957.50 $ 3,044.50 $ 87 Higher
Operating Profit(Loss) $ 10,010 $ 9,995 ($ 15)* Lower

* Please enter positive figure if table does not enter negative figure.

(2) Yes, as the net income has been increased.

(3) Minimum Quantity :

Sales Price of Special Order =$ 47

Variable Cost of Special Order = ($ 23 + 14.5 + 6.5) = $ 44

Additional Fixed Cost              = $ 87,000

Minimum Quantity                   = Additional Fixed Cost / CM per Unit

                                              = $ 87,000 / ($ 47 - $ 44)

= 29,000 Units

Working :

Status Quo 455,000 Units Alternativ 479,000 Units
Sales Revenue 455,000 * $ 65 = $ 29,575 (455,000 * $ 65) + (24,000 * $ 47) = $ 30,703
Less : Variable Costs
Direct Material 455,000 * $ 18.50 = $ 8,417.5 (455,000 * $ 18.50) + 24,000 * ($ 18.50 + $ 4.50) = $ 8,969.50
Direct Labor 455,000 * $ 12.50 = $ 5,687.5 (455,000 * $ 12.50) + 24,000 * ($ 12.50 + $ 2.00) = $ 6,035.5
Variable OH 455,000 * $ 5.50 = $ 2,502.5 (455,000 * $ 5.50) + 24,000 * ($ 5.50 + $ 1.00) = $ 2,658.50
Fixed OH $ 2,957.50 $ 2,957.50 + $ 87 = $ 3,044.50

Please vote up and write your doubts in comment section. Would be glad to help you further.